OpenStudy (anonymous):

Hi, Can you help me find the derivative of f(x)=2ln(7x)? Thanks.

OpenStudy (idku):

y=2ln(7x) and you need dy/dx.

OpenStudy (idku):

the derivative of y, is dy/dx. the derivative of ln(7x) is by the chain rule, the inner argument "7x" is going to be over one, like this : 1/(7x) and times the derivative of the inner function, times the derivative of "7x" (which is 7) So d/dx 2 ln (7x) = 2 (1 / 7x ) (7) = 14 / (7x) = 2/x So, dy/dx = 2/x

OpenStudy (anonymous):

Wow, this is very clear. Thank you!